# Learning from History: Task-agnostic Model Contrastive Learning for Image Restoration

### 18 Dec 2023 v2

@@ Abstract
Contrastive learning has emerged as a prevailing paradigm for high-level vision tasks, which, by introducing properly negative samples, has also been exploited for low-level vision tasks to achieve a compact optimization space to account for their ill-posed nature. However, existing methods rely on manually predefined and task-oriented negatives, which often exhibit pronounced task-specific biases. To address this challenge, our paper introduces an innovative method termed ’learning from history’, which dynamically generates negative samples from the target model itself. Our approach, named Model Contrastive paradigm for Image Restoration (MCIR), rejuvenates latency models as negative models, making it compatible with diverse image restoration tasks. We propose the Self-Prior guided Negative loss (SPN) to enable it. This approach significantly enhances existing models when retrained with the proposed model contrastive paradigm. The results show significant improvements in image restoration across various tasks and architectures. For example, models retrained with SPN outperform the original FFANet and DehazeFormer by 3.41 dB and 0.57 dB on the RESIDE indoor dataset for image dehazing. Similarly, they achieve notable improvements of 0.47 dB on SPA-Data over IDT for image deraining and 0.12 dB on Manga109 for a 4$\times$ scale super-resolution over lightweight SwinIR, respectively. Code and retrained models are available at https://github.com/Aitical/MCIR.
@@ Introduction
Image restoration, aiming at recovering a high-quality image from the degraded one, is a fundamental problem in the fields of image processing and computer vision (Yang et al. 2021; Wang, Chen, and Hoi 2021a; Jiang et al. 2023). Deep learning approaches have brought about considerable advancements in image restoration while there are still challenges due to its ill-posed nature (Wang, Chen, and Hoi 2021b; Lu et al. 2023). The success of self-supervised learning paradigm for high-level tasks, especially those contrastive learning methods, has drawn great attention (Gui et al. 2023a). This inspires many researchers to made strides in improving the end-to-end learning paradigm for image restoration tasks, incorporating the concept of sample contrastive learning and bridging the gap between high-level and low-level tasks (Wu et al. 2021; Chen et al. 2022b; Ye et al. 2022; Wu, Jiang, and Liu 2023; Zheng et al. 2023). Image restoration tasks usually contain high-quality ground truth as learning target (positive sample), and more attention is paid on mining appropriate negative examples. For instance, Wu et al. (Wu et al. 2021) directly utilized the low-quality input as the negative sample and introduced the contrastive paradigm for the image dehazing task. Wu et al. proposed a hard negative construction for image super-resolution tasks. Compared to the super-resolved results, hard negatives with similar image quality to the anchor sample can push it ahead effectively (Wu, Jiang, and Liu 2023). Most recently, Zheng et al. proposed a progressively improved negative lower bound for image dehazing, employing multiple pre-trained models to provide consensual negatives (Zheng et al. 2023).
@@
While existing contrastive learning methods for low-level tasks, leveraging potent approaches like hard negative mining (Wu, Jiang, and Liu 2023) and curriculum learning strategies (Zheng et al. 2023), have shown impressive performance, certain intrinsic limitations persist. Chiefly, there exists an over-reliance on task-oriented prior, leading to limited generalization capability across multiple image restoration tasks. Many existing methods exhibit a pronounced task-oriented bias, where negative sample generation is often influenced by prior knowledge and empirical evaluation centered around a singular target task (Wu, Jiang, and Liu 2023; Zheng et al. 2023). Such approaches, while effective in isolation, often impede to a diverse set of image restoration tasks and model architectures. In light of these challenges, it compels us to ask: Is there a task-agnostic and general method for negative sampling that could potentially enhance the performance of a diverse range of image restoration tasks?
@@
Taking this into consideration, we turn our focus to the target model itself, rather than sample selection. In recent literature, much of the spotlight has been directed towards negative sample collection, often overlooking a latent gem—the latency model—during the learning process. Specifically, the latency model, when operating within a close optimization step, shares strikingly similar parameters with the current model. This intrinsic similarity paves the way for the construction of proper negatives pertinent to the current anchor sample. To illustrate this concept, we introduce a toy example in Fig. 1. It becomes evident that, throughout the learning journey, the output of the latency model exhibits a sub-optimal but congruent distribution relative to the current anchor sample. This alignment offers the potential to derive ’hard’ negatives that are well-suited to the task at hand. Furthermore, as the entire learning process is incrementally refined, the negatives in our model contrastive paradigm adopt a curriculum way naturally. Motivated by these insights, we put forth an innovative model contrastive paradigm for image restoration tasks. In contrast to the previous approaches with complex negative mining strategies, the core of our proposed model contrastive paradigm lies in the self-promoted negatives between latency and current models. Notably, it is task-agonist and general to diverse image restoration tasks.
@@
In detail, we proposed a model contrastive paradigm for various image restoration tasks (MCIR) with a Self-Prior guided Negative loss (SPN). Importantly, the proposed MCIR is compatible with existing approaches across different architectures and tasks. We retrain existing models by the proposed model contrastive and some results are presented in Fig. 2. One can find that retrained models achieve promising improvement across multiple tasks and architectures. More specifically, based on the lightweight EDSR (Lim et al. 2017) and SwiniR (Liang et al. 2021), our retrained models display superior performance on Manga109 test dataset for ×$\times 4$ scale image super resolution, boasting improvements of 0.16 dB and 0.12 dB in term of PSNR. Retrained IDT (Xiao et al. 2022) achieves 0.38 dB and 0.7 dB gains on Rain200L and SPA datasets for image deraining. Moreover, for image dehazing, we gain a notable improvement of 3.41 dB and 0.57 dB compared to the original FFANet (Qin et al. 2020a) and DehazeFormer (Song et al. 2023), respectively.
@@
  • The paper proposes a novel approach MCIR through a task-agnostic model contrastive paradigm, which provides the adaptive generation of negative samples directly from the target model itself. Unlike conventional methods that manually apply negative samples to specific target task, the proposed model contrastive paradigm exhibits versatility across multiple tasks and models.
  • The paper introduces the Self-Prior guided Negative loss (SPN) for image restoration, seamlessly compatible with existing methods. SPN provides a simple to enhance existing image restoration models by integrating self-supervision principles within our model contrastive paradigm.
@@
Figure 2: Comparisons between models retrained by our proposed model contrastive paradigm and the originals. Retrained models can achieve remarkable improvements on various image restoration tasks. Details are provided in Section Experiment
@@
Image restoration, aiming at recovering a high-quality image from the degraded one, is a fundamental problem in the fields of image processing and computer vision. It is indispensable in several domains, including photography, surveillance, medical imaging, and remote sensing (Su, Xu, and Yin 2022). In recent years, approaches based on deep learning have emerged as potent tools, providing significant advances for image restoration (Wang, Chen, and Hoi 2021a; Jiang et al. 2023; Wang et al. 2021a; Gui et al. 2023b). Among these, Convolutional Neural Networks (CNN) have been utilized, resulting in a plethora of methods designed to address various image restoration tasks (Dong et al. 2016; Zhang et al. 2017; Zhang, Zuo, and Zhang 2018; Yang et al. 2017a). Furthermore, remarkable strides have been made by the development of a spectrum of deep models, including proposals for residual or dense architectures (Lim et al. 2017), the integration of attention mechanisms (Zhang et al. 2018; Liu et al. 2019), and the implementation of UNet-based architectures (Ronneberger, Fischer, and Brox 2015; Jiang et al. 2020). In addition, some work explored the attention mechanism to boost the performance for image super-resolution (Zhang et al. 2018; Liu et al. 2018; Mei, Fan, and Zhou 2021), image deraining (Li et al. 2018a; Wang et al. 2019a), image deblurring (Suin, Purohit, and Rajagopalan 2020), and so on. On the frontier of recent development, Transformer architectures achieves significant breakthroughs in computer vision (Liu et al. 2021; Dosovitskiy et al. 2021). This trend has spurred the development of Transformer-based interventions for image restoration tasks such as image super-resolution (Liang et al. 2021), image dehazing (Song et al. 2023; Zamir et al. 2022), image deraining (Xiao et al. 2022), image deblurring (Wang et al. 2022), and low-light image enhancement (Wang et al. 2023). We are witnessing a surge in the performance of increasingly complex architectures and there is a growing demand for applications in real-world systems. Many researcher have turned attention to lightweight methods (Jiang et al. 2021; Lamba and Mitra 2021; Hui et al. 2019; Wu et al. 2023b; Lu et al. 2022; Wu et al. 2023a).
@@
Self-supervised learning has achieved great success in high-level vision tasks, especially those contrastive learning methods (Gui et al. 2023a). For low-level image restoration tasks, many researchers attempted to bridge the gap between low-level and high-level tasks and exploit more self-supervised information to promote image restoration tasks (Wang et al. 2021b; Wu et al. 2021; Chen et al. 2022b; Ye et al. 2022; Zheng et al. 2023; Wu, Jiang, and Liu 2023). In contrast to unsupervised learning for high-level vision tasks, low-level tasks usually have high-quality ground-truth images. Discovering informative negative samples is crucial for low-level tasks. Wu et al. took the low-quality input as the negative sample and proposed a perceptual-based contrastive loss to achieve a more compact convergence space (Wu et al. 2021). Wu et al. proposed a practical contrastive learning framework for super-resolution single image and proposed the hard negative construction (Wu, Jiang, and Liu 2023). By interpolating the negative information, they further improve the performance for target image restoration task. However, adaptation and generalization of these methods to various image restoration models and tasks are limited.
Compared to existing methods (Wu et al. 2021; Wu, Jiang, and Liu 2023; Zheng et al. 2023), our work presents a task-agnostic model contrastive paradigm for image restoration, utilizing a self-prior guided negative loss. This approach simplifies negative sample generation and offers a versatile framework for various image restoration tasks, setting it apart from existing methods.
@@
Figure 3: Comparisons of different learning paradigms for image restoration tasks. (a) Depicts the vanilla end-to-end learning paradigm, where target image restoration models are trained with supervised reconstruction loss; (b) Illustrates the sample contrastive paradigm adopted in existing methods, leveraging additional information from manually defined negative samples; (c) Showcases our model contrastive paradigm, which generates negative samples from the latency target model, providing a task-agnostic and adaptive approach to various image restoration tasks.
@@
Deep learning-based image restoration methodologies have recently made considerable breakthroughs across multiple restoration tasks, including image super-resolution, image dehazing, image deblurring, image deraining, and so on. In this paper, we focus our attention on the vanilla end-to-end learning process in deep learning methods. To facilitate clarity, we provide a succinct illustration of the end-to-end image restoration framework, visualized in Fig. 3 (a). Given a low-quality input image denoted as $I^{LQ}$, the target model $f_{\theta}$, where $\theta$ represents model parameters, processes low-quality input to produce a reconstructed image, $I^{Rec}$. Concurrently, its high-quality counterpart is denoted by $I^{HQ}$ and the optimizing of the target model is steered by the minimization of the reconstruction loss $\mathcal{L}_{rec}$. Delving further, Fig. 3 (b) showcases the contrastive learning strategies tailored for image restoration tasks. Typically, researchers harness the power of negative samples, exploiting informative task-prior specific negatives and extend the vanilla end-to-end learning paradigm by the negative constraint $\mathcal{L}_{neg}$ (Wu et al. 2021; Zheng et al. 2023; Wu, Jiang, and Liu 2023).
In this paper, we propose an innovative model contrastive paradigm for image restoration tasks, shown in Fig. 3 (c). Instead of utilizing task-specific and predefined negative samples, we craft our negatives directly from the target model itself. At this point, we denote the target model $f_{\theta}$, indicative of the optimized one at the current training iteration $t$. Subsequently, we employ a distinct step, $s$, to establish a latency model $f_{\theta^{{}^{\prime}}}$, also referred to as the negative model, to generate adaptive negative samples for the training of the target model. The overall pipeline is straightforward, and the detailed implementation of negative models and associated loss functions will be discussed in the following sections.
@@
We propose a practical implementation for leveraging latency models to generate informative negative samples during training, as introduced earlier. This method circumvents the challenges of significant gaps between model checkpoints and the impracticality of frequent processing, especially with larger models. In detail, we introduce the efficient strategy of exponential moving averages (EMA) to achieve a smooth negative model. The update equation for the negative model is:
$\begin{split}&\theta^{{}^{\prime}}=w\theta^{{}^{\prime}}+(1-w)\theta,\\ &\mathrm{s.t.}~{}t\%s=0.\end{split}$ (1)
In this schema, $w$ denotes the update weight, $s$ is the update step, and $t$ captures the current iteration. To ensure the preservation of latency parameters, we adopt a long step $s$ and selectively update the negative model $f_{\theta}$ at intervals of every $s$ iterations.
@@
At the heart of our model contrastive paradigm, a loss function mediates between the target reconstruction $I^{Rec}$ and its negative counterpart, $I^{Neg}$. Herein, we take the pre-trained VGG (Simonyan and Zisserman 2015) network as the embedding network to map the samples into a latent feature space, where $f^{Rec}=\text{VGG}(I^{Rec})$ and $f^{Neg}=\text{VGG}(I^{Neg})$. Then the proposed negative loss $\mathcal{L}_{neg}$ is formulated as
@@
where $N$ denotes the total number of negative models and $f_{i}^{Neg}$ corresponds to the latent feature of the $i$th negative sample.
Compared to existing methods, our negative regularization is a self-prior guided loss function, where the negative samples stem from the target model itself. More important, it is general and transferable to existing image restoration models while retaining the original learning strategy. Typically, image restoration tasks utilize a reconstruction loss, $\mathcal{L}{rec}$, relying on metrics such as Mean Absolute Error (MAE) and Mean Squared Error (MSE). To validate the prowess of our model-based contrastive paradigm by $\mathcal{L}_{neg}$, we have retrained numerous exiting methods with it, testing across different image restoration tasks and architectures. The precise formulation of the reconstruction loss $\mathcal{L}_{rec}$ is contingent on the method under retraining.
@@
Method Architecture Scale
@@
EDSR-light CNN ×$\times 2$ 32.06/0.9303 33.57/0.9175 32.16/0.8994 31.98/0.9272 30.54/0.9769
+SPN (Ours) 32.19/0.9313 33.67/0.9182 32.21/0.9001 32.23/0.9297 30.64/0.9772
@@
For image deblurring, we retrain and evaluate NAFNet (Chen et al. 2022a) on GoPro dataset.
@@
For image deraining, we conduct experiments on several publicly available benchmarks, namely Rain200L/H (Yang et al. 2017b), DID-Data (Li et al. 2018b), DDN-Data (Fu et al. 2017), and SPA-Data (Wang et al. 2019b). We adopt the Transformer-based IDT model (Xiao et al. 2022) as our baseline and subsequently retrain it using the proposed model contrastive paradigm.

### 19 Jan 2024 v4

@@ Abstract
Contrastive learning has emerged as a prevailing paradigm for high-level vision tasks, which, by introducing properly negative samples, has also been exploited for low-level vision tasks to achieve a compact optimization space to account for their ill-posed nature. However, existing methods rely on manually predefined and task-oriented negatives, which often exhibit pronounced task-specific biases. To address this challenge, our paper introduces an innovative method termed ’learning from history’, which dynamically generates negative samples from the target model itself. Our approach, named Model Contrastive paradigm for Image Restoration (MCIR), rejuvenates latency models as negative models, making it compatible with diverse image restoration tasks. We propose the Self-Prior guided Negative loss (SPN) to enable it. This approach significantly enhances existing models when retrained with the proposed model contrastive paradigm. The results show significant improvements in image restoration across various tasks and architectures. For example, models retrained with SPN outperform the original FFANet and DehazeFormer by 3.41 and 0.57 dB on the RESIDE indoor dataset for image dehazing. Similarly, they achieve notable improvements of 0.47 dB on SPA-Data over IDT for image deraining and 0.12 dB on Manga109 for a 4$\times$ scale super-resolution over lightweight SwinIR, respectively. Code and retrained models are available at https://github.com/Aitical/MCIR.
@@ Introduction
Image restoration, aiming at recovering a high-quality image from the degraded one, is a fundamental problem in the fields of image processing and computer vision (Yang et al. 2021; Wang, Chen, and Hoi 2021a; Jiang et al. 2023). Deep learning approaches have made considerable advancements in image restoration, while there are still challenges due to its ill-posed nature (Wang, Chen, and Hoi 2021b; Lu et al. 2023). The success of the self-supervised learning paradigm for high-level tasks, especially those using contrastive learning methods, has drawn great attention (Gui et al. 2023a). This inspires many researchers to make strides in improving the end-to-end learning paradigm for image restoration tasks, incorporating the concept of sample contrastive learning and bridging the gap between high-level and low-level tasks (Wu et al. 2021; Chen et al. 2022b; Ye et al. 2022; Wu, Jiang, and Liu 2023; Zheng et al. 2023). Image restoration tasks usually contain high-quality ground truth as the learning target (positive sample), and more attention is paid to obtaining appropriate negative examples. For example, Wu et al. (Wu et al. 2021) directly utilized low-quality input as the negative sample and introduced the contrastive paradigm for the image dehazing task. Wu et al. proposed a hard negative construction for image super-resolution tasks. Compared to the super-resolved results, hard negatives with similar image quality to the anchor sample can push it ahead effectively (Wu, Jiang, and Liu 2023). Most recently, Zheng et al. utilized multiple pre-trained models to provide consensual negatives and proposed a progressively improved negative lower bound for image dehazing (Zheng et al. 2023).
@@
While existing contrastive learning methods for low-level tasks, leveraging potent approaches such as hard negative mining (Wu, Jiang, and Liu 2023) and curriculum learning strategies (Zheng et al. 2023), have shown impressive performance, certain intrinsic limitations persist. Chiefly, there exists an over-reliance on task-oriented prior, leading to limited generalization capability across multiple image restoration tasks. Many existing methods exhibit a pronounced task-oriented bias, where negative sample generation is often influenced by prior knowledge and empirical evaluation centered on a singular target task (Wu et al. 2021; Wu, Jiang, and Liu 2023; Zheng et al. 2023). Such approaches, while effective in isolation, often impede to a diverse set of image restoration tasks and model architectures. In light of these challenges, it compels us to ask: Is there a task-agnostic and general method for negative sampling that could potentially enhance the performance of a diverse range of image restoration tasks?
@@
Taking this into consideration, we turn our focus to the target model itself, rather than sample selection. In recent literature, much of the spotlight has been directed towards negative sample collection, often overlooking a latent gem, the latency model, during the learning process. Specifically, the latency model, when operating within a close optimization step, shares strikingly similar parameters with the current model. This intrinsic similarity paves the way for the construction of proper negatives pertinent to the current anchor sample. To illustrate this concept, we introduce a toy example in Fig. 1. It becomes evident that, throughout the learning journey, the output of the latency model exhibits a suboptimal but congruent distribution relative to the current anchor sample. This alignment offers the potential to derive ’hard’ negatives that are well suited to the task at hand. Furthermore, as the entire learning process is incrementally refined, the negatives in our model contrastive paradigm adopt a curriculum way naturally. Motivated by these insights, we put forth an innovative model contrastive paradigm for image restoration tasks. In contrast to previous approaches with complex negative mining strategies, the core of our proposed model contrastive paradigm lies in the self-promoted negatives between latency and current models. Notably, it is a task-agonist and general to diverse image restoration tasks.
@@
In detail, we propose a model contrastive paradigm for various image restoration tasks (MCIR) with a Self-Prior guided Negative loss (SPN). Importantly, the proposed MCIR is compatible with existing approaches across different architectures and tasks. We retrain existing models by the proposed model contrastive and some results are presented in Fig. 2. Retrained models can be found to achieve promising improvement across multiple tasks and architectures. More specifically, based on lightweight EDSR (Lim et al. 2017) and SwinIR (Liang et al. 2021), our retrained models show superior performance on the Manga109 test dataset for ×$\times 4$ scale image super resolution, boasting improvements of 0.16 dB and 0.12 dB in terms of PSNR. Retrained IDT (Xiao et al. 2022) achieves 0.38 dB and 0.7 dB gains on Rain200L and SPA datasets for image deraining. Furthermore, for image dehazing, we gain a notable improvement of 3.41 dB and 0.57 dB compared to original FFANet (Qin et al. 2020a) and DehazeFormer (Song et al. 2023), respectively.
@@
  • The paper proposes a novel approach MCIR through a task-agnostic model contrastive paradigm, which provides the adaptive generation of negative samples directly from the target model itself. Unlike conventional methods that manually apply negative samples to a specific target task, the proposed model contrastive paradigm exhibits versatility across multiple tasks and models.
  • The paper introduces the self-prior guided negative loss (SPN) for image restoration, which is seamlessly compatible with existing methods. SPN provides a simple to enhance existing image restoration models by integrating self-supervision principles within our model contrastive paradigm.
@@
Figure 2: Comparisons between models retrained by our proposed model contrastive paradigm and the originals. Retrained models can achieve remarkable improvements on various image restoration tasks.
@@
Image restoration, crucial in image processing and computer vision, aims to restore high-quality images from degraded versions (Su, Xu, and Yin 2022). Deep learning, particularly CNNs, has revolutionized this domain (Wang, Chen, and Hoi 2021a; Jiang et al. 2023; Wang et al. 2021a; Gui et al. 2023b), with advancements in various CNN-based methods (Dong et al. 2016; Zhang et al. 2017; Zhang, Zuo, and Zhang 2018; Yang et al. 2017a), and the incorporation of novel architectures such as residual networks (Lim et al. 2017), attention mechanisms (Zhang et al. 2018; Liu et al. 2019), and UNet structures (Ronneberger, Fischer, and Brox 2015; Jiang et al. 2020). As well as some lightweight architectures were proposed for practical application (Jiang et al. 2021; Lamba and Mitra 2021; Hui et al. 2019; Wu et al. 2023b; Lu et al. 2022; Wu et al. 2023a). Recent trends include Transformer-based architectures (Liu et al. 2021; Dosovitskiy et al. 2021), delivering breakthroughs in many image restoration tasks (Liang et al. 2021; Song et al. 2023; Zamir et al. 2022; Xiao et al. 2022; Wang et al. 2022, 2023). In this work, we turn our attention to the optimizing strategy and exploit an effective model contrastive learning method to refresh the performance existing models.
@@
Self-supervised learning, particularly contrastive methods, has seen remarkable success in high-level vision tasks (Gui et al. 2023a). In the field of low-level image restoration, researchers are exploring the integration of self-supervised regularization to bridge the gap between low- and high-level tasks (Wang et al. 2021b; Wu et al. 2021; Chen et al. 2022b; Ye et al. 2022; Zheng et al. 2023; Wu, Jiang, and Liu 2023). Typically, low-level tasks often have high-quality ground truth, making the identification of informative negative samples vital. Wu et al. employed low-quality inputs as negative samples and introduced a perceptual-based contrastive loss for more effective convergence (Wu et al. 2021). Wu et al. proposed a practical contrastive learning framework for single image super-resolution tasks, enhancing performance through hard negative construction and negative information interpolation (Wu, Jiang, and Liu 2023). Nonetheless, the adaptability and generalization of these methods across diverse image restoration models and tasks remain a challenge.
Our work distinguishes itself from existing methods (Wu et al. 2021; Wu, Jiang, and Liu 2023; Zheng et al. 2023) by introducing a task-agnostic model contrastive paradigm for image restoration. This novel approach shifts the focus from negative samples to negative models, offering a straightforward yet effective framework for a variety of image restoration tasks.
@@
Deep learning-based image restoration methodologies have recently made considerable breakthroughs across multiple restoration tasks, including image super-resolution, image dehazing, image deblurring, image deraining, and so on. In this paper, we focus our attention on the vanilla end-to-end framework with deep learning methods.
Given a low-quality input image $I^{LQ}$, the target model $f_{\theta}$ (where $\theta$ denotes the model parameters) processes this input to produce a reconstructed image $I^{Rec}$. The high-quality counterpart of the image is represented as $I^{HQ}$. The optimization of $f_{\theta}$ is guided by minimizing the reconstruction loss $\mathcal{L}_{rec}$, defined as:
$\min_{\theta}\mathcal{L}_{rec}(f_{\theta}(I^{LQ}),I^{HQ}),$ (1)
where $\mathcal{L}_{rec}$ typically depends on metrics such as Mean Absolute Error (MAE) and Mean Squared Error (MSE).
Delving further, existing contrastive learning methods for image restoration (Wu et al. 2021; Zheng et al. 2023; Wu, Jiang, and Liu 2023) utilize negative samples $I^{Neg}$, extending the standard end-to-end learning with a negative loss $\mathcal{L}_{neg}$:
$\min_{\theta}\mathcal{L}_{rec}(f_{\theta}(I^{LQ}),I^{HQ})+\mathcal{L}_{neg}(f_{\theta}(I^{LQ}),I^{Neg}),$ (2)
where the negative samples are predefined and generated based on task-specific priors, which limit their generalization capability across diverse image restoration tasks.
In this work, we introduce a novel model contrastive paradigm for image restoration, shifting from task-specific negative samples to negatives derived from the target model itself. We designate the target model as $f_{\theta}$, representing its state in a given training iteration $t$. A key innovation is the development of a latency model, $f_{\theta^{{}^{\prime}}}$ (the negative model), which is utilized to produce adaptive negative samples for training $f_{\theta}$. This approach simplifies the overall process in existing methods and provides a general framework for image restoration. The specifics of implementing negative models and the corresponding loss functions will be elaborated in subsequent sections.
@@
We propose a practical implementation for leveraging latency models to generate informative negative samples during training, as introduced earlier. This method avoids the challenges of significant gaps between model checkpoints and the impracticality of frequent processing, especially with larger models. In detail, we introduce the efficient strategy of exponential moving averages (EMA) to achieve a smooth negative model. The update equation for the negative model is as follows:
$\theta^{{}^{\prime}}=w\theta^{{}^{\prime}}+(1-w)\theta,~{}\mathrm{s.t.}~{}t\%s=0.$ (3)
In this schema, $w$ denotes the update weight, $s$ is the update step, and $t$ captures the current iteration. To ensure the preservation of latency parameters, we adopt a long step $s$ and selectively update the negative model $f_{\theta}$ at intervals of every $s$ iteration.
@@
At the heart of our model contrastive paradigm, a loss function mediates between the target reconstruction $I^{Rec}$ and its negative counterpart, $I^{Neg}$. Here, we take the pre-trained VGG (Simonyan and Zisserman 2015) network as the embedding network to map the samples into a latent feature space, where $f^{Rec}=\text{VGG}(I^{Rec})$ and $f^{Neg}=\text{VGG}(I^{Neg})$. Then the proposed negative loss $\mathcal{L}_{neg}$ is formulated as
@@
where $N$ denotes the total number of negative models and $f_{i}^{Neg}$ corresponds to the latent feature of the $i$-th negative sample.
Compared to existing methods, our negative regularization is a self-prior guided loss function, where the negative samples stem from the target model itself. More important, it is general and transferable to existing image restoration models while retaining the original learning strategy. Typically, image restoration tasks utilize a reconstruction loss, $\mathcal{L}{rec}$, relying on metrics such as Mean Absolute Error (MAE) and Mean Squared Error (MSE). To validate the prowess of our model-based contrastive paradigm by $\mathcal{L}_{neg}$, we have retrained numerous exiting methods with it, testing across different image restoration tasks and architectures. The precise formulation of the reconstruction loss $\mathcal{L}_{rec}$ is dependent on the retraining method.
@@
Methods Architecture Scale
@@
EDSR-light CNN ×$\times 2$ 34.06/0.9303 33.57/0.9175 32.16/0.8994 31.98/0.9272 38.54/0.9769
+SPN (Ours) 34.19/0.9313 33.67/0.9182 32.21/0.9001 32.23/0.9297 38.64/0.9772
@@
In this section, we apply the following four image restoration tasks to test and validate the effectiveness of the proposed method.
@@
For image deblurring, we retrain and evaluate NAFNet (Chen et al. 2022a) on GoPro dataset (Nah, Hyun Kim, and Mu Lee 2017).
@@
For image deraining, we conduct experiments on several publicly available benchmarks, namely Rain200L/H (Yang et al. 2017b), DID-Data (Li et al. 2018), DDN-Data (Fu et al. 2017), and SPA-Data (Wang et al. 2019). We adopt the Transformer-based IDT model (Xiao et al. 2022) as our baseline and subsequently retrain it using the proposed model contrastive paradigm.